(m = million/s ; b = billion/s)
Conifers Ltd, a tree relocation company, follows a stable plus special dividend policy. The company recently (2020) had earnings available to common shareholders of R500m on sales of R2b and has 250m shares outstanding. The company maintains a stable pay-out ratio of 5% and issues special dividends whenever its net profit margin exceeds 10%. Special dividends amount to all profits exceeding 50% of the amount after the stable dividend had been deducted.
Required:
Determine the total dividend that Conifers Ltd will pay per share in 2020.
1
Expert's answer
2020-09-28T19:13:23-0400

"DPS=\frac{D-SD}{S}"

"D=500\times0.05=25"

"SP=(500-25)\times0.5\times0.05=11.875"

"DPS=\frac{D-SD}{S}=\frac{25-11.875}{250}=0.0525"